OpenStudy (anonymous):

find the integral of 1/sqrt(x^2-9)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{x ^{2}-9} }\] right?

OpenStudy (anonymous):

is that the question?

OpenStudy (sirm3d):

use the substitution \[\large x = 3\sec \theta\]